OpenStudy (anonymous):

x^3 -2x^2+13x=0 leave in exact form no decimals approx

OpenStudy (anonymous):

You could write the same equation as: x(x^2 - 2x + 13) = 0 so one root is x = 0 and the others are the roots of x^2 - 2x + 13 =0 equation which are (1 - 2sqrt3) i and (1 + 2sqrt3 i)

OpenStudy (anonymous):

did you solved this using quadratic formula negative b + or - square root b^2 - 4ac over 2a?

OpenStudy (radar):

I believe he did and he simplified the answer.

OpenStudy (anonymous):

i tried that i did broke it down i got 2+ or - sqrt-48 over 2 times "a" which is 2 but when you divide the top (2+ or - sqrt-48) i got 24 which breaks down to to 2i sqrt 6 how did he get the 3?

OpenStudy (radar):

\[\sqrt{-48}=\sqrt{16\times-3}\]

OpenStudy (radar):

\[\sqrt{16}\times \sqrt{-3}\]\[4\sqrt{-3}\]

OpenStudy (radar):

does that simplification rings a bell?

OpenStudy (anonymous):

yeah i get it now its hard when someone just leaps from one spot to another

OpenStudy (radar):

I haven't double checked but nikola usually gets it right. Good luck

OpenStudy (anonymous):

Sweets, http://en.wikipedia.org/wiki/Quadratic_equation a=1 b=(-2) c=13 b^2 - 4ac = (-2)^2 - 4*1*13 = -48 which is smaller than 0,so roots are complex and can be calculated as : -b/2a + (sqrt [- (b^2 - 4ac)] /2a)*i -b/2a - (sqrt [- (b^2 - 4ac)] /2a)*i

OpenStudy (anonymous):

yay nikola lol

OpenStudy (radar):

And I have to agree with you sometimes simplified looks more complicated lol\[\sqrt{-48}=4\sqrt{-3}\] doesn't look to much more simple Hi

OpenStudy (anonymous):

2+(sqrt[-2^2-4(1)(13)]/2(1) = 2+(sqrt[-48])/2 = so do I divide out the 2 from the bottom like so or no lol sorry for all this... 1+(sqrt[-24] then you break down the 24 to 4i sqrt 6

OpenStudy (anonymous):

so 1+\[4i \sqrt{6} and 1-4i \sqrt{6}\]

OpenStudy (anonymous):

that was suppose to be 2i sqrt6 not the 4 my bad

OpenStudy (anonymous):

Sweets, a=1 b=(-2) c = 13 b^2 - 4ac = (-2)^2 - 4*1*13 = -48 < 0 so; -b/2a + (sqrt [- (b^2 - 4ac)] /2a)*i -b = -(-2) = 2 so -b/2a is 2/2*1 =1 sqrt [- [b^2 - 4ac) ] = sqrt [-(-48)] = sqrt48 which is 4sqrt3 (sqrt [- (b^2 - 4ac)] /2a)*i is 4sqrt3 / 2 *1 = 2sqrt3 I hope this helps.

OpenStudy (anonymous):

i have got 2+24i and 2-24i as the answer

OpenStudy (anonymous):

yes thats correct nikola:)

OpenStudy (anonymous):

Sweets,it s -b/2*a not -b/a :)

OpenStudy (radar):

\[(-(-2)\pm \sqrt{(-2)^{2}-(4)(1)(13)})/2\]\[(2\pm \sqrt{4-52})/2\]\[(2\pm \sqrt{-48})/2\] can you take it from there?

OpenStudy (anonymous):

Oooooo lol i see i see

OpenStudy (anonymous):

you miss one thing it messes it all up lol

OpenStudy (anonymous):

yeah...thats what math is all about:)

OpenStudy (anonymous):

haha gotta love it

OpenStudy (anonymous):

so th eanswer probably is 1+24i & 1-24i if i am not wrong:)

OpenStudy (radar):

Continuing on.\[(2\pm(\sqrt{16\times-3})/2\] \[(2\pm(\sqrt{16}\sqrt{-3})/2\]\[(2\pm(4\sqrt{-3})/2\] now do the division\[1\pm2\sqrt{-3}\]

OpenStudy (anonymous):

how did u get -3 there? under the root???

OpenStudy (radar):

I would leave it there are you could go further and do this. Convert the radical as follows:\[\sqrt{-3}=\sqrt{3 X-1}\]

OpenStudy (radar):

The square root of -1 is the imaginary operator i so the final answer becomes:\[1\pm \sqrt{3}i\]

OpenStudy (radar):

Did you understand I was trying to show in the radical 3 times a -1 ? and i left out the 2 so it should be\[1\pm2\sqrt{3}i\]

OpenStudy (anonymous):

yeah i seen what you did 4 squared times 3 is 48 but you had to divide the 4 by 2 i get it

OpenStudy (radar):

Yeahh!!!Now practice is the key. Notice that they did not want you to really get the value no decimals etc. so that is as far as you need to take it.

OpenStudy (anonymous):

lol thank all of you so much. Im not as slow as i seem this was just a total brain fart

OpenStudy (anonymous):

Im a visual learner so it helped when you showed :)

OpenStudy (anonymous):

you're welcome sweets,good luck with your studies.